This is a question from the Circles and Arcs subject of my Geometry book. I am clueless about what formula to use. Please help, the problem is: "The wheel of one car has a diameter of 20 inches. The wheel of another car has a diameter of 24 inches. About how many more revolutions must the smaller wheel make than the larger wheel to travel 100 feet?"
Answer by shree840(260)

2*pi*r is the circumeference or pi*d
using pi as 3.14
first car 3.14*20=62.8 or 62.8/12=5.23 ft
to cover 100 ft needs 100/5.23 rev=19.12
2nd car pi*d= 3.14*24inches=3.14*24/12=6.28
rev needed foer 100 ft= 100/6.28=15.92 diff 19.12-15.92=3.19 ANS